OpenStudy (anonymous):

can someone help me with this? he didn't lecture on finding points other than when t=0, so i am very, very lost. thanks. don't know where to start. (1 pt) Use a graph to estimate the coordinates of the highest point and the leftmost point on the curve x=4te^t,y=4te^−t. Then find the exact coordinates. Highest point: . Leftmost point: . The curve has one horizontal asymptote and one vertical asymptote. Find them. Horizontal asymptote: y= . Vertical asymptote: x= .

OpenStudy (anonymous):

@amistre64 , you've helped before. can you help me now?

OpenStudy (anonymous):

ok, well, i have to go to bed. i'll check back later and hopefully someone will have answered me by then...

OpenStudy (amistre64):

im thinking like this: the leftmost point is when x stops moving and turns back around x = 4t e^t x' = 4 e^t + 4t e^t = 4e^t (1+t) since 4e^t is always positive .... at t = -1 the slope is 0 at t < -1, the slope is negative at t > -1 the slope is positive